OpenStudy (anonymous):

What is the value of x? −x = 6 − 4x + 18

OpenStudy (anonymous):

I think it is -8

OpenStudy (anonymous):

Close, but it would be x=8.

OpenStudy (anonymous):

I think you just got a negative mixed up somewhere.

OpenStudy (anonymous):

Let's solve your equation step-by-step. −x=6−4x+18 Step 1: Simplify both sides of the equation. −x=6−4x+18 Simplify: (Show steps) −x=−4x+24 Step 2: Add 4x to both sides. −x+4x=−4x+24+4x 3x=24 Step 3: Divide both sides by 3. 3x/3=24/3 x=8 Answer: x=8

OpenStudy (anonymous):

here you go Simplifying 2x + 6 = 4x + -2 Reorder the terms: 6 + 2x = 4x + -2 Reorder the terms: 6 + 2x = -2 + 4x Solving 6 + 2x = -2 + 4x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-4x' to each side of the equation. 6 + 2x + -4x = -2 + 4x + -4x Combine like terms: 2x + -4x = -2x 6 + -2x = -2 + 4x + -4x Combine like terms: 4x + -4x = 0 6 + -2x = -2 + 0 6 + -2x = -2 Add '-6' to each side of the equation. 6 + -6 + -2x = -2 + -6 Combine like terms: 6 + -6 = 0 0 + -2x = -2 + -6 -2x = -2 + -6 Combine like terms: -2 + -6 = -8 -2x = -8 Divide each side by '-2'. x = 4 Simplifying x = 4

OpenStudy (anonymous):

x=8 not x=-8